1. In the reaction Mg(OH)2 + 2HCl → MgCl2 + 2H2 O, you were given 15


grams of Mg(OH)2 and 25 grams of HCl. Which of the two reactants is the


limiting reagent? How much MgCl2 will be produced in the reaction? (Mg(OH)2


= 58.32 g/mol, HCl = 36.46 g/mol, MgCl2 = 95.32 g/mol)


1
Expert's answer
2022-04-29T11:08:04-0400

From the equation 1 mole of Mg(OH)2 reacts with 2 moles of HCl.

That means that "58.32 g" of Mg(OH)2 react with "2*36.46=72.92 g" of HCl.

15 g of Mg(OH)2 will need "15*72.92\/58.32 = 18.76g" of HCl. Since we have 25 g of HCl, Mg(OH)2 is the limitting reagent.

Find the mass of MgCl2:

"m(MgCl_2) = m(Mg(OH)_2)*M(MgCl_2)\/M(Mg(OH)_2) = 15*95.32\/58.32 = 24.52 (g)"

Answer: Mg(OH)2 is the limitting reagent. 24.52 g of MgCl2 will be produced in the reaction.